1) Calculate the number of C atoms in 0.334 mole of C. Nc=______atoms of C

2) Calculate the number of SO2 molecules in 2.73 moles of SO2. NSO2=______ molecules of SO2

3) Calculate the moles of Fe in 2.82x1022 atoms of Fe. Nfe=_______ mole of Fe

4) Calculate the moles of C2H6O in 3.11x1024 molecules of C2H6O. nC2H6O=_______ moles of  C2H6O
Answer #1

1)

use:

number of atoms = number of moles * Avogadro’s number

= 0.334 * 6.022*10^23 atoms

= 2.011*10^23 atoms

Answer: 2.01*10^23 atoms

2)

use:

number of molecules = number of mol * Avogadro’s number

number of molecules = 2.73 * 6.022*10^23 molecules

number of molecules = 1.644*10^24 molecules

Answer: 1.64*10^24 molecules

3)

use:

number of mol = number of atoms / Avogadro’s number

number of mol = 2.82*10^22 / 6.022*10^23 mol

number of mol = 4.683*10^-2 mol

Answer: 4.68*10^-2 mol

4)

use:

number of mol = number of molecules / Avogadro’s number

number of mol = 3.11*10^24 / 6.022*10^23 mol

number of mol = 5.164 mol

Answer: 5.16 mol